# Spectral gap of a convex combination of a random permutation and a deterministic matrix

### v3

@@ Abstract
We study the spectral gap behavior of an operator obtained by summing a random permutation $M$ and a deterministic bistochastic matrix $Q$. We are interested in the asymptotic in terms of dimension. In the case where $(M,Q)$ are asymptotically free with amalgamation over the diagonal, we can compute limit operators $(u,q)$ which give the weak limit spectral distribution. Therefore we introduce free with amalgamation operators that are suitable for computing the spectral gap limit of our operator in high dimensions. We then approximate the spectral radius of the corresponding limit operator and finally give an upper bound for the spectral radius of the finite-dimensional operator. In particular, we show that if the deterministic matrix underlying graph is an expander, then the underlying graph associated to the sum with a random permutation is again an expander.
@@ 1.1 Notations and conventions
[…]
For a bounded operator $b\in\mathcal{B}\left(\mathrm{H}\right)$ where $\mathrm{H}$ is an Hilbert space, we denote by $\|b\|$ its operator norm associated to the inner scalar product norm of $\mathrm{H}$ and we denote by $b^{\ast}$ the adjoint operator of $b$. For $\epsilon\in\{+,-\}$, we set $b^{\epsilon}=b$ if $\epsilon=+$ and $b^{\epsilon}=b^{\ast}$ if $\epsilon=-$. Regarding the spectrum, for $B$ an operator on a Hilbert space of dimension $N$, we consider the eigenvalues with multiplicities. We denote the $k$-th eigenvalue by $\lambda_{k}(B)$ and order them as follows
@@
We call spectral gap of $B$ the following positive function of $B$
$\gamma(B):=\left|\lambda_{1}(B)\right|-\left|\lambda_{2}(B)\right|.$
@@
$\|B\|_{0\rightarrow 1}$ $:=\underset{x\in[N]}{\max}|\{y\in[N]\leavevmode\nobreak\ ;\leavevmode\nobreak\ B_{xy}\neq 0\}|$
$\|B\|_{0\rightarrow 1}^{\ast}$ $:=\max\{\|B\|_{0\rightarrow 1},\|B^{\ast}\|_{0\rightarrow 1}\}.$
Besides we introduce the norm
$\|B\|_{\infty\to\infty}^{\ast}:=\max\{\|B\|_{\infty\to\infty},\|B^{\ast}\|_{\infty\to\infty}\}$
where $\|\cdot\|_{\infty\to\infty}$ is the subordinate norm to the infinite norm on $\mathbb{C}^{N}$ i.e.
$\|B\|_{\infty\to\infty}:=\max_{\|X\|_{\infty}=1}\|BX\|_{\infty}=\max_{x}\sum_{y}|B_{xy}|.$
@@
$\kappa_{L}(B):=\underset{\begin{subarray}{c}k\leq p\leq L,\\ (x_{t},y_{t})\in[N]^{2k}\end{subarray}}{\max}\leavevmode\nobreak\ \left\{\max_{\begin{subarray}{c}n_{1}+\cdots+n_{k}=p\\ n_{t}\leq\ell(B)\end{subarray}}\leavevmode\nobreak\ \prod_{t=1}^{k}(B^{n_{t}})_{x_{t}y_{t}}\right\}^{\frac{1}{p+k}},$ (1)
where $\ell(B):=\frac{\ln(N)}{7\ln(\|B\|_{0\rightarrow 1}^{\ast})}$ and we set
@@ 1.2 Model
    $Q\boldsymbol{1}=\delta\boldsymbol{1}\quad\text{ and }\quad Q^{\ast}\boldsymbol{1}=\delta\boldsymbol{1}$
@@
    There exists $10<\alpha<$ 1 such that we have
    $d:=\underset{x\in[N]}{\max}|\{y\in[N]\leavevmode\nobreak\ ;\leavevmode\nobreak\ Q_{xy}\neq 0\}|\vee\underset{x\in[N]}{\max}|\{y\in[N]\leavevmode\nobreak\ ;\leavevmode\nobreak\ Q^{\ast}_{xy}\neq 0\}|=e^{o\left(\ln(N)^{\alpha}\right)}.$
We see $Q$ as an adjacency matrix for a $\delta$-regular graph. We then take interest in the matrix
$A=M+Q$ (3)
@@
where $M$ is a uniformly distributed permutation. In the case where $\delta=1$, $Q$ is called doubly stochastic. The set of bistochastic matrices being a convex polytope, any convex combination of a permutation matrix and $Q$ remains bistochastic. In this paper we also are interested in the following matrix
$A_{r}:=(1-r)M+rQ$ (4)
@@
where $M$ is a uniformly distributed permutation and $10<r<$ 1. If we consider $Q$ and $M$ as transition matrices of Markov chains on the same vertex set $[N]$, $P$ is the sum of these two graphs, rescaled to obtain a Markov chain. According to Perron-Frobenius Theorem, $1$ (resp. $2$) is then the largest eigenvalue of $A_{r}$ (resp. of $A$). We are interested in the asymptotic behavior in dimension $N$ of the spectral gap $1-|\lambda_{2}(A_{r})|$ (resp. $2-\left|\lambda_{2}(A)\right|$). The main result of this article is a probabilistic upper bounds on $|\lambda_{2}(A)|$ and $\left|\lambda_{2}(A_{r})\right|$ that depends only on parameters of $Q$. More precisely the bound we are interested in is the spectral radius of an operator that depends on $Q$ and $r$ and is independent with the permutation matrix $M$.
1.3 Free probability and spectral analysis of random matrices
One way of expressing the interest of free probability in terms of spectral analysis for large random matrices is that it is a suitable framework for studying the asymptotic behaviour of non-commutative random variables. Roughly speaking, the notion of freeness in non-commutative probability is the non-commutative equivalent of independence for (classical commutative) random variables. When considering independent non-commutative variables, such as random matrices, this independence is indeed translated into freeness asymptotically. Voiculescu [Voi91] introduced free probability with the idea of constructing spaces of operators imitating basic probability theory and we refer to Anderson, Guionnet and Zeitouni [AGZ10] for a comprehensive study of random matrices.
The asymptotic freeness of a couple $(A,B)$ of random matrices allows the construction of an explicit operator $\mathrm{P}_{\mathrm{free}}\in\mathcal{B}(\mathrm{H})$ such that the spectrum (resp. the empirical spectral measure) of $\mathrm{P}$a polynomial in $(A,B)$, converges asymptotically to the spectrum (resp. spectral measure) of $\mathrm{P}_{\mathrm{free}}$ in a sense yet to be defined. For instance, considering independent uniformly distributed random permutations matrices, the asymptotic operators are the left representations of the free group. For any integer $d$ we then denote $\mathbf{u}=(u(s))_{s\in[d]}$ the generators of $\mathbb{F}_{d}$ the free group with $d$ generators. We keep the same notation $(u(s))_{s\in[d]}$ for the left-representations, i.e. the operators defined by
$u(s):\ell^{2}(\mathbb{F}_{d})$ $\rightarrow\ell^{2}(\mathbb{F}_{d})$
$\delta_{g}$ $\mapsto\delta_{u(s)g}$
where $\delta_{g}:=(\mathds{1}(h=g))_{h\in\mathbb{F}_{d}}\in\ell^{2}(\mathbb{F}_{d})$. The linear form $\tau:=\langle\cdot\delta_{e},\delta_{e}\rangle$ has to be seen as the non-commutative analog of the expectation $\mathbb{E}(\cdot)$ in the classic probability setting. Then two non-commutative operators ${a}_{1},{a_{2}}\in\mathcal{B}\left(\ell^{2}\left(\mathbb{F}_{d}\right)\right)$ such that $\langle{a}_{i}\delta_{e},\delta_{e}\rangle=0$ are called free if for any non-trivial word ${a}$ made from the alphabet $({a}_{1},{a}_{2})$ we have $\langle{a}\delta_{e},\delta_{e}\rangle=0$. This is the analog of defining the classic independence of two complex random variable. Seen as a family of operators, $\mathbf{u}$ is a $d$-Haar free family. Then the fact that a family of possibly random unit matrices $\mathbf{M}=(M(s))_{s\in[d]}$ are asymptotically free means that for any non-commutative polynomial $P(\cdot)\in\operatorname{\mathbb{C}}\langle X(s),X(s)^{\ast}|s\in[d]\rangle$ we have:
$\lim_{N\rightarrow\infty}\operatorname{Tr}_{N}[P(\mathbf{M})]=\langle P(\mathbf{u})\delta_{e},\delta_{e}\rangle$ (5)
where $e\in\mathbb{F}_{d}$ is the neutral element. In general, for any family of possibly random matrices $\mathbf{M}$ and any family of operators $\mathbf{u}$, if the convergence (5) occurs for all polynomial, we say that the family $\mathbf{M}$ converges in distribution to the family $\mathbf{u}$. If such a convergence occurs it already gives spectral information over some operators obtained with the family $\mathbf{M}$. If one considers $P(\cdot)$ a symmetric polynomial, $\mathrm{P}=P(\mathbf{M})$ and $\mathrm{b}=P(\mathbf{u})$ then the previous convergence implies:
where $\mu_{\mathrm{b}}$ is the spectral measure given by the spectral theorem applied to $\mathrm{b}$ and the convergence is weak. When the operator $\mathrm{b}$ is bounded, its spectral measure is characterized by its moments. If we furthermore assume freeness of the family $\mathbf{u}=(u(s))_{s\in[d]}$ then the moments, and therefore the spectral measure, of the operator $\mathrm{b}$ are computable knowing the spectral measure of $u(s)$ for all $s\in[d]$. For the simple case of the sum of two free non-commutative operators, the operation on measures is called free convolution and one can refer to [AGZ10, Section 5.3.3]. To see the interest of asymptotic freeness in our setting, recall that we take interest in a random matrix, given by either (3) or (4), of the form
$A=U+Q$
Theorem 1.1 (Voiculescu, 1991).
For instance it means that if $Q$ is in the setting of the previous theorem and if one considers the operator $P=\frac{1}{4}\left(Q+Q^{\ast}+U+U^{\ast}\right)$, then the asymptotic spectral measure is completely characterized by the asymptotic distribution of $Q$. The fact that we softened the regularity of the distribution for the random matrix $U$ by taking a random permutation instead of a Haar distributed unitary, weakens the freeness of the asymptotic operators. This new freeness is called freeness with amalgamation, introduced in Section 2. This new freeness has a particular interest in random graph theory because the previous theorem holds again replacing freeness with freeness with amalgamation and invariance under $\mathrm{G}_{N}$ conjugation by invariance under $\mathrm{S}_{N}$ conjugation.
Theorem 1.2.
[ACD+18, Theorem 1.3] If $Q\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ converges in distribution and is invariant under $\mathrm{S}_{N}$ conjugation and $M\in\mathrm{S}_{N}$ is uniformly distributed, then the couple $(Q,M)$ is asymptotically free with amalgamation.
Once again the convergence in distribution and the freeness with amalgamation of the limit operators allows to compute the distribution of the limit family. An explicit algorithm to compute the joint distribution of free with amalgamation operators is given by Belinschi, Mai, Speicher [BMS17]. The previous setting is better suited for graph theory applications, since we can compute random graphs which joint entry wise law is not invariant under $\mathrm{O}_{N}$ or $\mathrm{U}_{N}$ conjugation but under $\mathrm{S}_{N}$. Indeed one can consider $Q$ the adjacency matrix (resp. the Markov operator of the simple walk) of an Erdős-Rényi graph, then the couple $(Q,M)$ is asymptotically free with amalgamation and the asymptotic spectral distribution of the operator defined by (3) (resp. by (4)) can be computed. The previous theorem is exactly the reason why we will, in Section 4, construct the explicit free with amalgamation operators $(q,u)$, so that when $Q$ is invariant with conjugation under $\mathrm{S}_{N}$, then 2.2 also says that we have no outliers, that is to say eigenvalues outside of the spectral support of $(1-r)u+rq$ or $u+q$.
1.4 Application to expansion of graphs and stronger convergences
@@
A more local approach to the same question is to ask whether, for a given graph, there is a not too small subset of the vertex set that is poorly connected to the rest of the vertex set. In computer science this phenomenon is called a bottleneck. We introduce an important tool that quantifies the previous questions, the Cheeger constant of $G=(V,E)$ a finite connected undirected graph. It is defined by the following:
@@
where for $F\subset V$, $\partial F\subset E$ is the set of edges connecting a vertex of $F$ to a vertex of $V\backslash F$. The Cheeger constant of a graph is a common quantifier of graph expansion. For $G=(V,E)$ a graph and $\epsilon>0$ we say that $G$ is a $\epsilon$-expander if we have
@@
where $\pi$ is the invariant measure of $\mathrm{P}$ and for $\nu$ a measure on $[N]$, $\|\nu\|_{TV}:=\frac{1}{2}\sum_{x}\nu(x)$ denotes the total variation norm. The link between the Cheeger constant and the spectral analysis of either the adjacency operator or the Markov operator for the simple walk is given by the following inequality.
@@
Let $G=(V,E)$ be a connected, $d$-regular undirected graph. We denote by $\mathrm{P}$ the Markov operator of the simple walk on $G$. We have:
@@
A $d$-regular graph $G=(V,E)$ is a $\epsilon$-expander for eigenvalues if:
@@
To apply the previous convergences to expansion of possibly random graphs one can consider the simple walk defined by the following Markov operator:
$\mathrm{P}=\frac{1}{2d}\sum_{s=1}^{d}M(s)+M(s)^{\ast}$ (8)
where $\mathbf{M}=(M(s))_{s\in[d]}$ is a family of permutations or tensor products of permutation. The asymptotic operator $\mathrm{b}=\frac{1}{2d}\sum_{s}u(s)+u(s)^{\ast}$ is then the adjacency operator of $\mathbb{F}_{d}$ Cayley graph. The corresponding spectral measure is given by Kesten-McKay law $\mu_{\mathrm{b}}(\mathrm{d}x)=f(\cdot)\mathrm{d}x$ where the density is given by:
$f:\mathbb{R}$ $\rightarrow\operatorname{\mathbb{R}}$
$x$ $\mapsto\frac{d}{2\pi}\frac{\sqrt{4(d-1)-x^{2}}}{d^{2}-x^{2}}\mathds{1}(|x|\leq 2\sqrt{d-1}).$
Having asymptotic freeness for the family $\mathbf{M}$, i.e. the convergence given in (5) and therefore the convergence (6), also means:
$|\sigma(\mathrm{P})\cap[-2\frac{\sqrt{d-1}}{d},2\frac{\sqrt{d-1}}{d}]^{\mathrm{c}}|=o(N).$
In particular we have at most $o(N)$ outliers, i.e. eigenvalues greater in modulus than the Alon-Boppana bound $2\sqrt{d-1}/d$. Also we have as expected that the Alon-Boppana bound is sharp since the weak convergence gives also that for all $\epsilon>0$ we have $|\sigma(\mathrm{P})\cap[2\sqrt{d-1}/d-\epsilon,2\sqrt{d-1}/d]|>\delta$ for some $\delta>0$ and $N$ large enough. The asymptotic freeness (5) had been proved for $\mathbf{M}$ a family of Haar distributed unitaries of $\mathrm{U}_{N}$ by Voiculescu [Voi91] in 1991 and for random permutations by Nica [Nic93] in 1993. In the case where we consider $Q$ a deterministic hermitian matrix, invariant under $\mathrm{S}_{N}$ conjugation, for instance the adjacency operator of an Erdős-Rényi, and $M$ a uniformly radom permutation, then considering
$\mathrm{P}:=Q+M+M^{\ast}$
one can apply 1.2 and, denoting $(q,u)$ the limit free with amalgamation operators for the joint distribution of $(Q,M)$ we have again
$|\sigma(\mathrm{P})\cap[-\|b\|,\|b\|]^{\mathrm{c}}|=o(N)$
where $b=q+u+u^{\ast}$.
In any case, the convergence given in (5) is not sufficient to conclude about the expansion of operators. In fact, it is precisely the absence of outliers that is needed to satisfy the expander 1.4. One way to obtain these expansions is to show strong asymptotic freeness. Again sticking to the unitary case, we will say that a family of unitary matrices $\mathbf{M}=(M(s))_{s\in[d]}$ is strongly asymptotically free if for any non-commutative polynomial $P(\cdot)\in\operatorname{\mathbb{C}}\langle X(s),X(s)^{\ast}|s\in[d]\rangle$ we have (5) and furthermore
$\lim_{N\rightarrow\infty}\|P(\mathbf{M})\|=\|P(\mathbf{u})\|,$ (9)
where the norms above are the operator norms associated to the inner-product of the Hermitian spaces $\operatorname{\mathbb{C}}^{N}$ and $\ell^{2}(\mathbb{F}_{d})$. Bordenave and Collins [BC19] showed the following theorem.
Theorem 1.5 (Bordenave, Collins 2018).
Let $d\geq 1$ be an integer and $\mathbf{M}=(M(s))_{s\in[d]}\in\mathrm{S}_{N}$ iid uniformly distributed random permutations. The family $\mathbf{M}$ is strongly asymptotically free in probability.
The previous theorem says that the random $d$-regular graph which simple walk operator is given by (8) is an expander and optimal. Indeed for any undirected $d$-regular graph there exists a constant $c_{d}$ such that we have the following Alon-Boppana bound for its simple walk operator $\mathrm{P}$
$\lambda_{2}(\mathrm{P})\geq\frac{2\sqrt{d-1}}{d}-\frac{c_{d}}{\ln(N)^{2}}=\|b\|-\frac{c_{d}}{\ln(N)^{2}},$
see [Moh10]. If now $\mathrm{P}$ is taken as in (8) with random permutations, the previous theorem gives
$\mathbb{P}(\lambda_{2}(\mathrm{P})\geq\frac{2\sqrt{d-1}}{d}+\epsilon)\underset{N\rightarrow\infty}{\rightarrow}0$
for all $\epsilon>0$. The previous convergence is exactly what we want to show for the operator $P$ given by (4).
Remark 1.1.
In the general setting we consider in this paper, that is to say $Q$ doubly stochastic, non-Hermitian and $P$ defined by (4) with $M$ a random permutation, 1.2 does not imply the convergence (6). Indeed we still have asymptotic freeness with amalgamation, but the fact that we are dealing with non-Hermitian operators invalidates the spectral theorem for the operator $p=(1-r)u+rq$. One can then consider the Brown measure of the operator $p$, and ask if Girko’s method applies to show the convergence (6) (see for instance [BC12]). This convergence would then depends the distribution chosen for $Q$.
Finally, an interest of looking directly at the operator $A=Q+M$ is that its structure and the computations of its moments are closer to the more general case where we study $A^{\prime}=a_{0}\otimes\operatorname{Id}_{N}+a_{1}\otimes Q+a_{2}\otimes M$ with $a_{i}$ fixed matrices of dimension $k$. Indeed, in many cases the strong convergence in the free case is showed using a linearization trick saying that if the spectrum of $A^{\prime}$ for any choices of $(a_{i})_{i}$ such that $A^{\prime}$ is Hermitian converges then we also have convergence of the spectrum of $P(M,Q)$ for any symmetric polynomial $P(\cdot)$ (see for instance [BC19, Section 6]). To adapt this linearization trick to our case is the fact that the matrix $a_{1}\otimes Q$ does not verify (H:1).
2 Freeness with amalgamation and main result
2.1 Definitions
We consider $Q\in\mathrm{M}_{N}(\mathbb{C})$ verifying (H:1), (H:2) and (H:3) and $M$ as before a uniformly distributed random permutation. We search for deterministic operators, denoted $(u,q)$, such that an approximation of the spectral radius of the deterministic operators
$a=u+q\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ p=(1-r)u+rq$ (10)
asymptotically bounds the spectral radius of the matrices defined by (3) and (4). As we explained in Section 1.3 we want to construct the couple $(u,q)$ free with amalgamation and such that $q$ has same $\ast$-distribution than $Q$ (see 2.1 below). These notions can be found in the article by Au, Cébron, Dahlqvist, Gabriel and Male or [ACD+18] or in Voiculescu, Dykema, and Nica [VDN92].

### v4

@@ Abstract
We study the spectral gap behavior of an operator obtained by summing a random permutation $M$ and a deterministic bistochastic matrix $Q$. We are interested in the asymptotic in terms of dimension. In the case where $(M,Q)$ are asymptotically free with amalgamation over the diagonal, we can compute limit operators $(u,q)$ which give the weak limit spectral distribution. Therefore we introduce free with amalgamation operators that are suitable for computing the spectral gap limit of our operator in high dimensions. We then approximate the spectral radius of the corresponding limit operator and finally give an upper bound for the spectral radius of the finite-dimensional operator. In particular, we show that if the deterministic matrix underlying graph is an expander, then the underlying Markov chain associated to the sum with a random permutation is again an expander.
@@ 1.1 Notations and conventions
For all integer $N\geq 1$ we set $N]:=\{1,..,N\}$, $\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ the algebra of complex matrices of dimension $N$, $\mathrm{D}_{N}(\operatorname{\mathbb{C}})$ its subalgebra of diagonal matrices and $\mathrm{S}_{N}\subset\mathrm{M}_{N}(\operatorname{\mathbb{C}})$ the symmetric group. We denote $\Delta:\mathrm{M}_{N}(\operatorname{\mathbb{C}})\rightarrow\mathrm{D}_{N}(\operatorname{\mathbb{C}})$ the projection defined by $\Delta(B)=(\delta_{ij}b_{ij})_{i,j\in[N]}$ where $\delta_{ij}=\mathds{1}(i=j)$ and $B=(b_{ij})_{i,j\in[N]}\in\mathrm{M}_{N}(\operatorname{\mathbb{C}})$. We denote $\operatorname{Tr}(\cdot)$ the classic trace operation on matrices and $\operatorname{Tr}_{N}(\cdot):=\frac{1}{N}\operatorname{Tr}(\cdot)$ the normalized trace. We denote by $\left|\cdot\right|$ the cardinality of a set and the usual absolute value. We denote by $\mathbb{P}(\cdot)$ and $\mathbb{E}(\cdot)$ the probability and expectation under the uniform measure on $\mathrm{S}_{N}$. We denote $\mathrm{M}_{N}(L^{\infty})$ the algebra of complex random matrices of dimension $N$ with essentially bounded entries and $\mathrm{D}_{N}(L^{\infty})$ its subalgebra of diagonal matrices. If $M$ is a random permutation matrix taken uniformly we denote $M\sim\mathcal{U}\left(\mathrm{S}_{N}\right)$.
We denote $\mathrm{B}(\mathrm{H})$ the algebra of bounded operator on the Hilbert space $\mathrm{H}$. For a bounded operator $b\in\mathrm{B}\left(\mathrm{H}\right)$ where $\mathrm{H}$ is an Hilbert space, we denote by $\|b\|$ its operator norm associated to the inner scalar product norm of $\mathrm{H}$ and we denote by $b^{\ast}$ the adjoint operator of $b$. For $\epsilon\in\{+,-\}$, we set $b^{\epsilon}=b$ if $\epsilon=+$ and $b^{\epsilon}=b^{\ast}$ if $\epsilon=$-. Regarding the spectrum, for $B$ an operator on a Hilbert space of dimension $N$, we consider the eigenvalues with multiplicities. We denote the $k$-th eigenvalue by $\lambda_{k}(B)$ and order them as follows
@@
$d=d(B)$ $:=\underset{x\in[N]}{\max}|\{y\in[N]\penalty 10000\ ;\penalty 10000\ B_{xy}\neq 0\}|\vee\underset{y\in[N]}{\max}|\{x\in[N]\penalty 10000\ ;\penalty 10000\ B_{xy}\neq 0\}|.$
@@
$\kappa_{L}(B):=\underset{\begin{subarray}{c}k\leq p\leq L,\\ (x_{t},y_{t})\in[N]^{2k}\end{subarray}}{\max}\penalty 10000\ \left\{\max_{\begin{subarray}{c}n_{1}+\cdots+n_{k}=p\\ n_{t}\leq\ell(B)\end{subarray}}\penalty 10000\ \prod_{t=1}^{k}(B^{n_{t}})_{x_{t}y_{t}}\right\}^{\frac{1}{p+k}},$ (1)
where $\ell(B):=\frac{\ln(N)}{7\ln(d(B)}$ and we set
@@ 1.2 Model
[…]
@@
    There exists $0<\alpha<1$ such that we have
    $d:=\underset{x\in[N]}{\max}|\{y\in[N]\penalty 10000\ ;\penalty 10000\ Q_{xy}\neq 0\}|\vee\underset{x\in[N]}{\max}|\{y\in[N]\penalty 10000\ ;\penalty 10000\ Q^{\ast}_{xy}\neq 0\}|=e^{o\left(\ln(N)^{\alpha}\right)}.$
We denote
$Q_{\perp}:=Q-\frac{1}{N}\boldsymbol{1}^{\top}\boldsymbol{1}.$ (3)
From (H:1) we have that $Q$ and $Q_{\perp}$ have a triangular representation in a common basis. We see $Q$ as an adjacency matrix for a $\delta$-regular graph. We then take interest in the operators, denoted $A_{\mathrm{x}}$ for $\mathrm{x}\in\{\mathrm{s},\mathrm{c}\}$. The first operator is given by the following sum
$A_{\mathrm{s}}=M+Q$ (4)
@@
where $M\sim\mathcal{U}(\mathrm{S}_{N})$. In the case where $\delta=1$, $Q$ is called doubly stochastic. The set of bistochastic matrices being a convex polytope, any convex combination of a permutation matrix and $Q$ remains bistochastic. In this paper we also are interested in the following convex combination
$A_{\mathrm{c}}:=(1-r)M+rQ$ (5)
@@
where $M\sim\mathcal{U}(\mathrm{S}_{N})$ and $0<r<1$. If we consider $Q$ and $M$ as transition matrices of Markov chains on the same vertex set $[N]$, $P$ is the sum of these two graphs, rescaled to obtain a Markov chain. According to Perron-Frobenius Theorem, $\lambda_{1}(A_{\mathrm{c}})=1$ (resp. $\lambda_{1}(A_{\mathrm{s}})=2$) is then the largest eigenvalue of $A_{\mathrm{c}}$ (resp. of $A_{\mathrm{s}}$). We are interested in the asymptotic behavior in dimension $N$ of the spectral gap $\lambda_{1}(A_{\mathrm{x}})-|\lambda_{2}(A_{\mathrm{x}})|$ for $\ast\in\{\mathrm{s},\mathrm{c}\}$. The main result of this article (see ˜1.4 and ˜2.1) is a probabilistic upper bounds on $|\lambda_{2}(A_{\mathrm{x}})|$ that depends only on parameters of $Q$. More precisely the bound we are interested in is the spectral radius of an operator that depends on $Q$ and $r$ and is independent with the permutation matrix $M$.
1.3 Graph expansion
@@
In this paper we study the second largest eigenvalue of the operator defined by (4) that can be seen as an adjacency matrix if $Q$ is an adjacency matrix and the operator given by (5) that can be seen as a Markov operator if $Q$ is a Markov operator. The motivation behind the main result of this article is to prove or improve the expansion of the graph or chain defined by $Q$ by adding the random edges defined by $M$.
To explain the link with the second largest eigenvalue of the operators $A_{\mathrm{s}}$ and $A_{\mathrm{c}}$, we first consider a more local approach of expansion. We ask whether, for a given graph, there is a not too small subset of the vertex set that is poorly connected to the rest of the vertex set. In computer science this phenomenon is called a bottleneck. The Cheeger constant of $G=(V,E)$ a finite connected undirected graph is a common quantifier of graph expansion and it is defined by the following
$h(G):=\min\{\frac{|\partial F|}{|F|};\penalty 10000\ F\subset V,\penalty 10000\ 0<|F|\leq\frac{1}{2}|V|\}$ (6)
@@
where $\pi$ is the invariant measure of $\mathrm{P}$ and for $\nu$ a measure on $[N]$, $\|\nu\|_{TV}:=\frac{1}{2}\sum_{x}\nu(x)$ denotes the total variation norm. The link between the Cheeger constant and spectral analysis is given by the following inequality.
@@
Let $G=(V,E)$ be a connected, $d$-regular undirected graph. We denote by $\mathrm{P}$ the Markov operator of the simple walk on $G$. We have
@@
A $d$-regular graph $G=(V,E)$ is a $\epsilon$-expander for eigenvalues if
@@
Even though we take interest in this article in non-Hermitian operators, for the reader understanding we only gave definitions for undirected graphs in this current section. Nevertheless we can similarly define the Cheeger constant and link it to the spectral expansion of a directed graph (see for instance [CHU05, Theorem 5.1]).
1.4 Free probability and spectral analysis of random matrices
Now that we gave some motivation for the study of spectrum in graph analysis, we introduce a tool for asymptotic spectral analysis for possibly random graphs. Roughly speaking, the notion of freeness in non-commutative probability is the non-commutative equivalent of independence for (classical commutative) random variables and therefore it is the suitable framework to study the asymptotic behavior of matrices spectrum. When considering independent non-commutative variables, such as random matrices, this independence is indeed translated into freeness asymptotically. Voiculescu [VOI91] introduced free probability with the idea of constructing spaces of operators imitating basic probability theory and we refer to Anderson, Guionnet and Zeitouni [AGZ10] for a comprehensive study of random matrices.
Freeness and convergence in distribution.
We have convergence in distribution for a couple $\mathbf{Q}=(Q_{1},Q_{2})$ of random matrices if for all $P$ non-commutative symmetric polynomial the sequence $\left(\operatorname{Tr}_{N}(P(\mathbf{Q},\mathbf{Q}^{\ast})\right)_{N\in\operatorname{\mathbb{N}}}$ converges. Assuming some control over the operator norms of the family $\mathbf{Q}$ this convergence implies the existence of a family $\mathbf{q}=(q_{1},q_{2})\in\mathrm{B}(\mathrm{H})^{2}$ of self-adjoint operators on a Hilbert space such that for all polynomial $P$ we have $\lim_{N\rightarrow\infty}\operatorname{Tr}_{N}(P(\mathbf{Q},\mathbf{Q}^{\ast}))=\langle P(\mathbf{q},\mathbf{q}^{\ast})\xi,\xi\rangle$ where $\xi\in\mathrm{H}$ verifies $\|\xi\|_{2}=1$. The linear form $\tau:=\langle\cdot\xi,\xi\rangle:\mathrm{B}(\mathrm{H})\rightarrow\operatorname{\mathbb{C}}$ has to be seen as the non-commutative analog of the expectation $\mathbb{E}(\cdot)$ in the classic probability setting. For the previous property of the convergence in distribution of random matrices one can refer to [AGZ10, Corollary 5.2.16]. In term of spectral analysis, considering $P(\cdot)$ a non-commutative symmetric polynomial we can apply the spectral theorem to the operator $\mathrm{b}=P(\mathbf{q},\mathbf{q}^{\ast})$ and denoting $\mathrm{P}=P(\mathbf{Q},\mathbf{Q}^{\ast})$ it means that there exists a unique probability measure $\mu_{\mathrm{b}}$ such that
where the previous convergence is weak. In particular if one can compute the distribution $\mu_{\mathrm{b}}$ and denoting $\rho(\mathrm{b})$ its spectral radius we have that for all $\delta>0$ there exists $\epsilon>0$ such that for all $N$ large enough
$\sigma(\mathrm{P})\cap[-\rho(\mathrm{b})+\delta,\rho(\mathrm{b})-\delta]\geq\epsilon N.$ (8)
In particular it gives the best expansion one can get for a graph with an adjacency matrix given by $\mathrm{P}$.
The freeness of the family $\mathbf{q}$ of operators (see [AGZ10, Section 5]) has to be seen as the analog of independence in classic probability. Freeness will translate asymptotically the independence and/or invariance of laws of random matrices and allow computability of the measure $\mu_{\mathrm{b}}$ (see [AGZ10, Section 5.3.3] and for the non-Hermitian case [BL99]). To see the interest of asymptotic freeness in our setting, recall that we take interest in a random matrix, given by either (4) or (5), of the form
$A=U+Q$
Theorem 1.2 (Voiculescu, 1991).
If we consider that $Q$ is in the setting of the previous theorem and $U$ is Haar distributed unitary independent of $Q$, then the asymptotic spectral measure of $P=\frac{1}{4}\left(Q+Q^{\ast}+U+U^{\ast}\right)$ is completely characterized by the asymptotic distribution of $Q$. The asymptotic freeness had been proved for a family of Haar distributed unitaries of $\mathrm{U}_{N}$ by Voiculescu [VOI91] in 1991 and for random permutations by Nica [NIC93] in 1993. In this article we want to consider random graphs, and the fact that we softened the regularity of the distribution for the random matrix $Q$ by considering random adjacency operators weakens the freeness of the asymptotic operators.
Freeness with amalgamation.
This new freeness is called freeness with amalgamation and we give here the definition. These notions can be found in the article by Au, Cébron, Dahlqvist, Gabriel and Male [ACD+18] or in Voiculescu, Dykema, and Nica [VDN92].